OpenStudy (anonymous):

If using calc. to input quadratic formula to solve 3x^2=-9/2x + 6 How would I key in?

OpenStudy (anonymous):

Hi Well, first put it in the standard form ax^2 + bx + c = 0 Then use the constants a, b, and c in the quadratic formula.

OpenStudy (anonymous):

So I have: a=3 b=9/2 c=-6

OpenStudy (anonymous):

looks good so far

OpenStudy (anonymous):

So, I can write it out OK, but if i want to key it into calc do I have to use \[\pm\] or parenthesis around the 9/2?

OpenStudy (anonymous):

usually the easiest thing to do is first compute what's in the radical, b^2 - 4ac. Take the square root. the +/- means you have two answers, one where you add that square root and one where you subtract it. Most calculators don't have a +/- key that does it automatically.

OpenStudy (anonymous):

If I were using a regular calculator, I'd probably first compute 4 * a * c, which is -72. Then key in 9/2, then square that, then add 72 (which is the same as subtracting -72). Then take the square root.

OpenStudy (anonymous):

I'm using TI-84, thought maybe I could plug it all together but that's just as easy

OpenStudy (anonymous):

I've never used one. I actually use mathematica personally, which solves the whole thing for you instantly anyway just by typing in the original equation.

OpenStudy (anonymous):

When I was in high school I used an old ti-35 which doesn't take complex expressions at all

OpenStudy (anonymous):

Yeah...I'm doing literally dozens of these quadratic formulas and it's starting to make my head hurt

OpenStudy (anonymous):

Thank you for the help

OpenStudy (anonymous):

THere are lots of quadratic calculators on the web, where you just type in A, B and C and it computes it for you

OpenStudy (anonymous):

For example, this one: http://www.math.com/students/calculators/source/quadratic.htm

OpenStudy (anonymous):

You can program it into your calculator as well, the language is pretty intuitive and easy to use, but you might have to google the syntax first.